A footballer kicks a ball at an angle of 45 degrees with the horizontal. If the ball was in the air for 10 seconds and lands 4000 meter away, determine its initial speed


1
Expert's answer
2020-03-13T11:04:38-0400

The horizontal distance for the projectile motion


"L=v_0t\cos\theta."

Hence, the initial speed


"v_0=\frac{L}{t\cos\theta}=\frac{4000}{10\cos45^{\circ}}=566\:\rm m/s."